\begin{thebibliography}{99}
\bibitem{Aspenberg2012}
Magnus Aspenberg and Walter Bergweiler,
Entire functions with Julia sets of positive measure.
Math. Ann. 352 (2012), 27--54.

\bibitem{Baker1968}
I.\ N.\ Baker,
Repulsive fixpoints of entire functions,
Math.\ Z. 104 (1968), 252-256.

\bibitem{Barrett2014}
Matthew Barrett and Alexandre Eremenko,
On the spherical derivative of a rational function.
Anal. Math. Phys. 4 (2014),
%no. 1-2,
73--81.

\bibitem{Bergweiler1993}
Walter Bergweiler,
Iteration of meromorphic functions.
Bull. Amer. Math. Soc. (N.S.) 29 (1993), 151--188.


\bibitem{Bergweiler1998}
Walter Bergweiler,
A new proof of the Ahlfors five islands theorem,
J.\ Analyse Math. 76 (1998), 337-347.

\bibitem{Bergweiler2005}
Walter Bergweiler,
Quasinormal families and periodic points.
In ``Complex analysis and dynamical systems II'',
Contemp. Math. 382, Amer. Math. Soc., Providence, RI, 2005,
pp. 55--63.

\bibitem{Bergweiler1995a}
Walter Bergweiler and Alexandre Eremenko,
On the singularities of the
inverse to a meromorphic function of finite order.
Rev. Mat.  Iberoamericana 11 (1995), 355--373.

\bibitem{Bergweiler2012}
Walter Bergweiler and Alexandre Eremenko,
On a property of the derivative of an entire function.
Ann. Acad. Sci. Fenn. Math. 37 (2012), 301--307.

\bibitem{Bergweiler1995} Walter Bergweiler and A. Hinkkanen,
On semiconjugation of entire functions.
Math. Proc. Cambridge Philos. Soc. 126 (1999), 565--574.

\bibitem{Bergweiler2009}
Walter Bergweiler, Bogus\l awa Karpi\'nska and Gwyneth M.\  Stallard, The growth
rate of an entire function and the Hausdorff dimension of its
Julia set.
J. London Math. Soc. (2) 80 (2009), 680--698.

\bibitem{Bergweiler2013}
Walter Bergweiler, Philip J. Rippon and Gwyneth M. Stallard,
Multiply connected wandering domains of entire functions.
Proc. London Math. Soc. (3) 107 (2013), 1261--1301.

\bibitem{Berteloot2010}
Fran\c ois Berteloot,
Lyapunov exponent of a rational map and multipliers of repelling cycles.
Riv. Math. Univ. Parma (N.S.) 1 (2010), 263--269.

\bibitem{Clunie1966}
J. Clunie and W. K. Hayman,
On the spherical derivative of integral and meromorphic functions.
Comment. Math. Helv. 40 (1966), 117--148.

\bibitem{Eremenko1989}
A. E. Eremenko,
On the iteration of entire functions.
In ``Dynamical systems and ergodic theory'',
Banach Center Publications 23, Polish Scientific Publishers, Warsaw, 1989, pp. 339--345.

\bibitem{Eremenko1990}
A. E. Eremenko and G. M. Levin,
Periodic points of polynomials.
Ukrainian Math. J. 41 (1989),
%no. 11,
1258--1262 (1990); translation of
Ukrain. Mat. Zh. 41 (1989),
%no. 11,
1467--1471, 1581.


\bibitem{Eremenko1992} A.\ E.\ Eremenko and M.\ Yu.\ Lyubich,
Dynamical properties of some classes of entire functions.
Ann.\ Inst.\ Fourier (Grenoble) 42 (1992), 989--1020.

\bibitem{Gelfert2010}
Katrin Gelfert, Feliks Przytycki and Micha{\l} Rams,
On the Lyapunov spectrum for rational maps.
Math. Ann. 348 (2010),
%no. 4,
965--1004.

\bibitem{Gelfert2013}
Katrin Gelfert, Feliks Przytycki, Micha{\l}  Rams and Juan Rivera-Letelier,
Lyapunov spectrum for exceptional rational maps,
Ann. Acad. Sci. Fenn. Math. 38 (2013), 631--656.

\bibitem{Goldberg2008}
Anatoly A. Goldberg and Iossif V. Ostrovskii,
Value distribution of meromorphic functions.
Amer. Math. Soc., Providence, RI, 2008.

\bibitem{Hayman1964}
W.\ K. Hayman,
Meromorphic functions.
Clarendon Press, Oxford, 1964.

\bibitem{Langley1995}
J.\ K.\ Langley, On the multiple points of certain meromorphic
functions. Proc. Amer. Math. Soc. 123 (1995), 355--373.

\bibitem{Levin2015}
Genadi Levin, Feliks Przytycki and Weixiao Shen,
The Lyapunov exponent of holomorphic maps.
Invent. Math.,
DOI 10.1007/s00222-015-0637-1

%\bibitem{Nevanlinna1953}
%Rolf Nevanlinna,
%Eindeutige analytische Funktionen.
%Springer, Berlin -- G\"ottingen -- Heidelberg, 1953.
%
\bibitem{Pommerenke1970}
Ch. Pommerenke,
Estimates for normal meromorphic functions,
Ann. Acad. Sci. Fenn., Ser. A I, no. 476 (1970), 10 pp.

\bibitem{Przytycki1994}
Feliks Przytycki, Letter to Alexandre Eremenko, \url{http://www.impan.pl/~feliksp/unpublished.html}, January 3, 1994.

\bibitem{Rippon2009}
P.\ J.\ Rippon and G.\ M.\ Stallard,
Functions of small growth with no unbounded wandering domains.
J. Anal. Math. 108 (2009), 61--86.

\bibitem{Rippon2011}
P.\ J.\ Rippon and G.\ M.\ Stallard,
Slow escaping points of meromorphic functions.
Trans. Amer. Math. Soc. 362 (2011), 4171--4201.

\bibitem{Rippon2012}
P.\ J.\ Rippon and G.\ M.\ Stallard,
Fast escaping points of entire functions.
Proc. London Math. Soc. (3) 105 (2012), 787--820.


\bibitem{Schleicher2010}
Dierk Schleicher,
Dynamics of entire functions.
In ``Holomorphic dynamical systems'',
Lecture Notes Math. 1998, Springer, Berlin, 2010, pp.~295--339.

\bibitem{Yao}
Xiao Yao, Daochun Sun and Zuxing Xuan,
A new geometric characterization of the Julia set.
Preprint, arXiv: 1512.05144.
\bibitem{Zdunik2014}
Anna Zdunik,
Characteristic exponents of rational functions.
Bull. Polish Acad. Sci. Math. 62 (2014), 257--263.
\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amsthm}
\usepackage{url}
\usepackage[showonlyrefs]{mathtools}
\mathtoolsset{showonlyrefs}
\numberwithin{equation}{section}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}{Lemma}[section]
\newtheorem{corollary}{Corollary}[section]
\theoremstyle{remark}
\newtheorem{remark}{Remark}[section]
\newtheorem*{ack}{Acknowledgment}
\def\R{\mathbb{R}}
\def\C{\mathbb{C}}

\def\N{\mathbb{N}}

\def\dist{\operatorname{dist}}
\def\re{\operatorname{Re}}
\def\im{\operatorname{Im}}
\def\Per{\operatorname{Per}}
\begin{document}
\title{Lyapunov exponents and related concepts for entire functions}
\author{Walter Bergweiler}
\address{Mathematisches Seminar,
Christian--Albrechts--Universit\"at zu Kiel,
Lude\-wig--Meyn--Str.~4, 24098 Kiel, Germany}
\email{bergweiler@math.uni-kiel.de}
\author{Xiao Yao}
\address{Department of Mathematical Sciences, Tsinghua University, P.\ R.\ China}
\email{yaoxiao0710107@163.com}
\author{Jianhua Zheng}
\address{Department of Mathematical Sciences, Tsinghua University, P.\ R.\ China}
\email{jzheng@math.tsinghua.edu.cn}
\subjclass[2010]{Primary 37F10; Secondary 30D05}
\thanks{The second author and the third author were supported by the grant (No. 11571193) of NSF of China}
\begin{abstract}
Let $f$ be an entire function and denote by $f^\#$ be the spherical derivative of~$f$
and by $f^n$ the $n$-th iterate of~$f$.
For an open set $U$ intersecting the Julia set $J(f)$, we consider how fast
$\sup_{z\in U} (f^n)^\#(z)$ and $\int_U (f^n)^\#(z)^2 dx\:dy$ tend to~$\infty$.
We also study the growth rate of the sequence $(f^n)^\#(z)$ for $z\in J(f)$.
\end{abstract}
\maketitle
\section{Introduction and results} \label{intro}
The Julia set $J(f)$ of a rational or entire function~$f$,
which we always assume to be neither constant nor rational of degree~$1$,
is the set of all points where the iterates $f^n$ of $f$ do not form a normal family.  Let
\begin{equation}\label{2a}
f^\#(z)=\frac{|f'(z)|}{1+|f(z)|^2}
\end{equation}
be the spherical derivative of~$f$.
Marty's theorem yields that a point $\xi\in\C$ is contained in $J(f)$ if and and only if
\begin{equation}\label{2a1}
\sup_{n\in\N}\sup_{z\in U} (f^n)^\#(z)=\infty
\end{equation}
for every neighborhood $U$ of $\xi$. Putting
\begin{equation}\label{17e}
\mu(U,f)=\sup_{z\in U} f^\#(z)
\end{equation}
we thus see that the sequence $(\mu(U,f^n))_{n\in\N}$  is unbounded.
It is not difficult to see that it actually tends to $\infty$ and
we are interested in the question how fast it tends to~$\infty$.

Let
\[
M(r,f) = \max_{|z|=r} |f(z)|
\]
be the maximum modulus of $f$ and
denote by $M^n(r,f)$ the iterate of $M(r,f)$ with respect to the first variable; that is,
\begin{equation}\label{2a2}
M^1(r,f)=M(r,f)
\quad\text{and}\quad
M^{n+1}(r,f)=M(M^n(r,f),f).
\end{equation}
It is easy to see that $M^n(R,f)\to\infty$ if $R$ is sufficiently large.
\begin{theorem}\label{thm7}
Let $f$ be an entire function and let $U$ be an open set intersecting the Julia
set of~$f$. Then, for any $R>0$, there exists $m\in \N$ such that
\begin{equation}\label{17f}
\mu(U,f^n)\geq \log M^{n-m}(R,f)
\end{equation}
for large~$n$.
\end{theorem}
For $f(z)=z^d$ we have $\log M^n(R,f)=d^n\log R$ and $\mu(U,f^n)\sim d^n/2$ as $n\to\infty$
if $U$ intersects the unit circle. So Theorem~\ref{thm7} gives the correct order of magnitude
for polynomials.

Next we show that analogous results hold if the supremum of the spherical derivative
is replaced by the normalized spherical area
\begin{equation}\label{2b}
S(U,f)=\frac{1}{\pi}\int_U f^\#(z)^2 dx\:dy.
\end{equation}
For a rational function $f$ of degree $d$ we have
\begin{equation}\label{2e}
c\, d^{n}\leq S(U,f^n)\leq d^{n}.
\end{equation}
for some positive constant $c$ and thus~\cite[Theorem~1]{Yao}
\begin{equation}\label{2e1}
\lim_{n\to\infty} \frac{1}{n}\log S(U,f^n)= \log d .
\end{equation}
Since $\mu(U,f)\geq \sqrt{S(U,f)}$ this implies that
\begin{equation}\label{2e3}
\lim_{n\to\infty} \frac{1}{n}\log \mu(U,f^n)\geq \frac12\log d .
\end{equation}
Barrett and Eremenko \cite[inequality (13) and the remarks following it]{Barrett2014}
showed that we always have strict inequality in~\eqref{2e3}, but that 
the constant $1/2$ on the right hand side
cannot be replaced by a larger constant. 

When dealing with rational functions, it is more systematical to consider
\begin{equation}\label{2c}
\|f'(z)\|=f^\#(z)(1+|z|^2)=|f'(z)|\frac{1+|z|^2}{1+|f(z)|^2}.
\end{equation}
instead of $f^\#(z)$, and this is the quantity considered in~\cite{Barrett2014}.
We note that Theorem~\ref{thm7} holds if $f^\#$ is replaced by
$\|f'\|$ in the definition of $\mu(U,f)$.  An analogous remark applies to the results below.

\begin{theorem}\label{thm5}
Let $f$ be a transcendental entire function and let $U$ be an open set intersecting the Julia
set of~$f$. Then, for any $R>0$, there exists $m\in \N$ such that
\begin{equation}\label{2l}
S(U,f^n)\geq \log M^{n-m}(R,f)
\end{equation}
for large $n$.
\end{theorem}
This result gives the right order of magnitude for the growth of $S(U,f^n)$.
\begin{theorem}\label{thm6}
Let $f$ be a transcendental entire function and let $U$ be a bounded open subset of~$\C$.
Then there exists $R>0$ such that
\begin{equation}\label{17b}
S(U,f^n)\leq \log M^{n}(R,f)
\end{equation}
for all $n\in\N$.
\end{theorem}
We note that it is easy to deduce from Theorems~\ref{thm7} and~\ref{thm5} that
\begin{equation}\label{2e2}
\lim_{n\to\infty} \frac{1}{n}\log \mu(U,f^n)=\infty
\quad\text{and}\quad
\lim_{n\to\infty} \frac{1}{n}\log S(U,f^n)=\infty .
\end{equation}
for a transcendental entire function~$f$.
The second equation answers a question from~\cite{Yao}, where it was shown that this
holds under various additional hypotheses.

We now consider how fast $(f^n)^\#(z)$ can tend to~$\infty$ for a point $z\in J(f)$.
A result of Przytycki says that for rational functions the maximal growth
rate of the sequence $((f^n)^\#(z))$ over all $z\in J(f)$ is essentially the same as the one obtained
when restricting to periodic points $z$ only.
More precisely, Przytycki showed (\cite{Przytycki1994}, the proof is reproduced in~\cite{Gelfert2013}) that
if $f$ is a rational function, then
\begin{equation}\label{2a9a}
\limsup_{n\to\infty}  \frac{1}{n} \sup_{z\in \C} \log  \|(f^n)'(z)\|
=
\sup_{z\in \Per(f)} \lim_{n\to\infty}  \frac{1}{n}\log  \|(f^n)'(z)\|,
\end{equation}
where $\Per(f)$ denotes the set of periodic points of~$f$.
Note that if $z$ is a periodic point of $f$, say $f^p(z)=z$ and $\lambda=(f^p)'(z)$, then
\begin{equation}\label{2a3}
\lim_{n\to\infty}\frac{1}{n}\log \|(f^n)'(z)\|=
\lim_{n\to\infty}\frac{1}{n}\log (f^n)^\#(z)
=\frac{\log|\lambda|}{p}.
\end{equation}
The limit on the right hand side is called the \emph{Lyapunov exponent} of $f$ at $z$ and denoted
by $\chi(f,z)$. More generally,
\begin{equation}\label{2a4}
\overline{\chi}(f,z)=\limsup_{n\to\infty}\frac{1}{n}\log (f^n)^\#(z)
\quad\text{and}\quad
\underline{\chi}(f,z)=\liminf_{n\to\infty}\frac{1}{n}\log (f^n)^\#(z)
\end{equation}
are called the \emph{upper} and \emph{lower Lyapunov exponent} of $f$ at~$z$; see, e.g., \cite{Gelfert2010,Gelfert2013,Levin2015}
for some recent results on Lyapunov exponents for rational maps.

On the left hand side of~\eqref{2a9a} one may replace the supremum over all $z\in\C$ by the supremum over all $z\in U$,
if $U$ is an open set intersecting $J(f)$. Thus~\eqref{2a9a} takes the form
\begin{equation}\label{2a9}
\lim_{n\to\infty} \frac{1}{n}\log \mu(U,f^n)
=
\sup_{z\in \Per(f)} \chi(f,z),
\end{equation}

Eremenko and Levin~\cite[Theorem~3]{Eremenko1990} showed that if $f$ is a polynomial of degree $d\geq 2$, then
there exists a periodic point $z$ such that $ \chi(f,z)\geq \log d$,
with strict inequality unless $f$ is conjugate to the monomial $z\mapsto z^d$.
It follows from~\eqref{2e3} and~\eqref{2a9a} that
if $f$ is a rational function of degree~$d\geq 2$, then
there exists a periodic point $z$ such that $ \chi(f,z)> (\log d)/2$;
see also \cite{Berteloot2010,Gelfert2013,Zdunik2014} for related results.
Finally, \eqref{2e2} and~\eqref{2a9} suggest that if $f$ is a transcendental entire function, then
\begin{equation}\label{2a7}
\sup_{z\in \Per(f)} \chi(f,z)=\infty.
\end{equation}
It follows from the results in~\cite{Bergweiler2005} that this is indeed the case.
\begin{theorem}\label{thm1}
Let $f$ be a transcendental entire function.  Then the set of all $z$
such that $\chi(f,z)=\infty$ is dense in $J(f)$.
\end{theorem}
The essential statement is here that there exists $z\in J(f)$ with $\chi(f,z)=\infty$.
Once this is known, it is easy to see that the set of all such points is dense in $J(f)$.
Note that such points cannot be periodic
since $\chi(f,z)<\infty$ for a periodic point $z$ by~\eqref{2a3}.

It seems plausible that Theorem~\ref{thm1} can be improved by giving a lower bound for
$(f^n)^\#(z)$ which depends on the maximum modulus of~$f$.
However, Theorem~\ref{thm4} below will show that such a lower bound will have to be much
smaller than that given in Theorems~\ref{thm7} and~\ref{thm5}.

We can give such a lower bound for functions in
the Eremenko-Lyubich class $B$ consisting of all transcendental entire functions for which
the set of critical and (finite) asymptotic values is bounded.
In fact, we only need to assume that $f$ has a logarithmic singularity over $\infty$.
This includes functions in $B$ since for such functions all
singularities over $\infty$ are logarithmic.

The lower order $\lambda(f)$ of an entire function $f$ is defined by
\begin{equation}\label{17a}
\lambda(f) = \liminf_{r \to \infty} \frac{\log \log M(r,f)}{\log r}.
\end{equation}
Taking the limes superior in~\eqref{17a} yields the order $\rho(f)$.
\begin{theorem}\label{thm3}
Let $f$ be a transcendental entire function with a logarithmic singularity over $\infty$.
Then the set of all $z$ such that
\begin{equation}\label{5p}
\liminf_{n\to\infty} \frac{1}{n} \log \log (f^n)^\#(z) \geq \log (1+ \lambda(f))
\end{equation}
is dense in $J(f)$.
\end{theorem}
If $f$ has a logarithmic singularity over $\infty$, then $\lambda(f)\geq 1/2$; see, e.g.,
\cite[Proof of Corollary~2]{Bergweiler1995a} or \cite[p.~1788]{Langley1995} for this observation.
Hence Theorem~\ref{thm3} yields the following result.
\begin{corollary}\label{thm2}
Let $f$ be a transcendental entire function with a logarithmic singularity over $\infty$.
Then the set of all $z$ such that
\begin{equation}\label{5p0}
\liminf_{n\to\infty} \frac{1}{n} \log \log (f^n)^\#(z) \geq \log\frac32
\end{equation}
is dense in $J(f)$.
\end{corollary}
Theorems~\ref{thm3} and Corollary~\ref{thm2} are sharp. More precisely, we have the
following result.
\begin{theorem}\label{thm4}
For each $\rho\in [1/2,\infty)$ there exists $f\in B$ with
$\lambda(f)=\rho(f)=\rho$ such that if $z\in \C$ satisfies $\chi(f,z)=\infty$, then
\begin{equation}\label{5p3}
\limsup_{n\to\infty} \frac{1}{n} \log \log (f^n)^\#(z) \leq \log (1+ \rho) .
\end{equation}
\end{theorem}

\begin{ack}
We thank Alexandre Eremenko and Lasse Rempe-Gillen for helpful comments.
\end{ack}

\section{Background from complex dynamics and function theory} \label{prelim}
For an introduction to the iteration theory of entire functions we refer to~\cite{Bergweiler1993,Schleicher2010}.
A basic result of the theory is the following lemma.
\begin{lemma}\label{rpp}
The Julia set of a transcendental entire function is the closure of the set of repelling periodic points.
\end{lemma}
For rational functions this result was obtained by both Fatou and Julia, for transcendental
entire functions it is due to Baker~\cite{Baker1968}.

The \emph{exceptional set} $E(f)$ of an entire function $f$ is the set of all
$z\in\C$ for which the backward orbit
\begin{equation} \label{8a}
O^-(z)=\bigcup_{n=0}^\infty f^{-n}(z)
\end{equation}
is finite. It is a simple consequence of Picard's theorem that $E(f)$ contains at most one point.
The following result is sometimes called the ``blowing-up property'' of the Julia set.
\begin{lemma}\label{excep-set}
Let $f$ be entire, $U\subset \C$ open with $U\cap J(f)\neq\emptyset$ and $K\subset\C\backslash E(f)$ compact.
Then $f^n(U)\supset K$ for all large $n\in\N$.
\end{lemma}

The \emph{escaping set}
\[
I(f) = \{ x \in \R^m \colon f^n(x) \to \infty \},
\]
introduced in~\cite{Eremenko1989}, plays an important role in transcendental dynamics. Its subset
\begin{equation}
\label{fe2}
A(f) = \left\{ z \in \C \colon \text{ there exists } l \in \N\text { with } |f^{n}(z)| > M(R,f^{n-l}) \text{ for } n\geq l\right\},
\end{equation}
where $R> \min _{z \in J(f)}|z|$ and $J(f)$ is the Julia set, is called the
\emph{fast escaping set}. It was introduced in~\cite{Bergweiler1995} and has also turned out
to be very useful in transcendental dynamics.
A thorough study of this set is given in~\cite{Rippon2012} where it is also shown that
\begin{equation}
\label{fe3}
A(f) = \left\{ z \in \C \colon \text{ there exists } l \in \N\text { with } |f^{n}(z)| > M^{n-l}(R,f) \text{ for } n\geq l\right\},
\end{equation}
with $R$ so large that $M^n(R,f)\to\infty$ as $n\to\infty$.
The equivalence of~\eqref{fe2} and~\eqref{fe3} is also apparent from the following lemma proved
in~\cite[Lemma 2.1]{Bergweiler2013}.
\begin{lemma}\label{lem-fe}
Let $f$ be a transcendental entire function and $\varepsilon>0$.
Then there exists $R> 0$ such that if $r > R$ and $n\in \N$, then
\[
 M((1+\varepsilon)r,f^n) \geq M^n(r,f).
\]
\end{lemma}
The following lemma (see, e.g., \cite[Lemma~2.2]{Rippon2009}) is a
a consequence of Hadamard's three circles theorem; that is, the convexity of $\log M(r,f)$ in $\log r$.
\begin{lemma}\label{lem-fe3}
Let $f$ be a transcendental entire function and $c>1$. Then
\[
\log M(r^c,f) \geq c\log M(r,f)
\]
for all  sufficiently large~$r$.
\end{lemma}
The next lemma can be found in~\cite{Eremenko1989} for the escaping  set and
in~\cite{Bergweiler1995,Rippon2012} for the fast escaping  set.
\begin{lemma}\label{JIA}
Let $f$ be entire. Then $J(f)=\partial I(f)=\partial A(f)$.
\end{lemma}
The next lemma consists of Koebe's distortion theorem and Koebe's one quarter theorem.
Here and in the following we denote by $D(a,r)$ the open disk of radius $r$ around~$a$.
\begin{lemma}\label{Koebe}
Let $g\colon D(a,r)\to \C$ be univalent, $0< \rho <1$ and $z\in D(a,\rho r)\backslash\{a\}$.
Then
\begin{equation}\label{ka}
\frac{1}{(1+\rho)^2}\leq \frac{|g(z)-g(a)|}{|g'(a)|\cdot |z-a|}\leq \frac{1}{(1-\rho)^2}
\end{equation}
and
\begin{equation}\label{kb}
\frac{1-\rho}{(1+\rho)^3}\leq \frac{|g'(z)|}{|g'(a)|} \leq \frac{1+\rho}{(1-\rho)^3}.
\end{equation}
Moreover,
\begin{equation}\label{kc}
g(D(a,r))\supset D\!\left(g(a),\frac{1}{4}|g'(a)|r\right).
\end{equation}
\end{lemma}
Koebe's  theorems are usually only stated  for  the special  case that $a=0$,
$r=1$, $g(0)=0$  and $g'(0)=1$, but the above  version follows
immediately from this special case.

The following lemma is Harnack's inequality.
\begin{lemma}\label{Harnack}
Let $u\colon D(a,r)\to \R$ be a positive harmonic function, $0< \rho <1$ and $z\in D(a,\rho r)$.
Then
\begin{equation}\label{harn}
\frac{1-\rho}{1+\rho}\leq \frac{u(z)}{u(a)} \leq \frac{1+\rho}{1-\rho}.
\end{equation}
\end{lemma}

\section{Proof of Theorem~\ref{thm7}} \label{proofthm7}
The following lemma is similar to results given in~\cite{Bergweiler2012,Clunie1966,Pommerenke1970}.
Here and in the following we denote by $D(a,r)$ the open disk around $a$ of radius~$r$.
\begin{lemma}\label{la1}
Let $f\colon D(a,r)\to\C$  be holomorphic and $K,L>0$.
Suppose that $|f(a)|\leq K$ and that $|f'(z)|\leq L$ whenever $|f(z)|= K$. Then
\begin{equation}\label{la1a}
|f(z)|< K \exp\!\left( \frac{2L}{K}|z-a|\right)
\quad\text{for}\ z\in D\!\left(a,\frac{r}{2}\right).
\end{equation}
\end{lemma}
\begin{proof}
We follow the arguments in~\cite[p.~303]{Bergweiler2012}
and  put $u(z)=\log (|f(z)|/K)$ so that $|\nabla u|=|f'/f|$.
With $G=\{z\in D(a,r) \colon |f(z)|>K\}$ we then have $a\notin G$ and
\begin{equation}\label{18a1}
|\nabla u(z)|\leq  \frac{L}{K}
\quad\text{for}\ z\in D(a,r)\cap \partial G.
\end{equation}
Let $z\in G\cap D(a,r/2)$ and put $d(z)=\dist(z,\partial G)$.
Since $a\notin G$ we have $d(z)\leq |z-a|$ and thus
there exists $z_1\in \partial G\cap \partial D(z,d(z))\cap D(a,r)$.
For $0<s<1$ we put $z_s=z+s(z_1-z)=sz_1+(1-s)z$ and deduce from Harnack's inequality that
\begin{equation}\label{18a}
u(z_s)\geq \frac{1-s}{1+s}u(z).
\end{equation}
It follows that
\begin{equation}\label{18b}
\frac{u(z)}{(1+s)d(z)}
\leq
\frac{u(z_s)}{(1-s)d(z)}
=
\frac{u(z_s)-u(z_1)}{(1-s)|z-z_1|}
=
\frac{u(z_s)-u(z_1)}{|z_s-z_1|}.
\end{equation}
Passing to the limit as $s\to 1$ we obtain
\begin{equation}\label{18c}
u(z)\leq 2|\nabla u(z_1)|d(z) \leq \frac{2L}{K} d(z) \leq  \frac{2L}{K}|z-a|
\quad\text{for}\ z\in  G\cap D\!\left(a,\frac{r}{2}\right),
\end{equation}
from which the conclusion follows.
\end{proof}
\begin{proof}[Proof of Theorem~\ref{thm7}]
Since, by Lemma~\ref{rpp}, repelling periodic points are dense in $J(f)$, we may assume
without loss of generality that $U=D(a,r)$ for some repelling periodic point $a$ and some $r>0$.
Since $a$ is periodic we have $|f^n(a)|\leq K$ for some $K$ and all~$n$.
Lemma~\ref{JIA} implies that there exists
$b\in A(f)\cap D(a,r/2)$.  Thus $|f^n(b)| \geq M^{n-m}(R,f)$ for some $m\in\N$ and all $n\geq m$.

With
\begin{equation}\label{17g}
L=\frac{K}{r} \log\frac{M^{n-m}(R,f)}{K}
\end{equation}
we thus have
\begin{equation}\label{17g1}
|f^n(b)|\geq K\exp\!\left(\frac{Lr}{K}\right)
> K\exp\!\left(\frac{2L}{K}|b-a|\right).
\end{equation}
Applying Lemma~\ref{la1} we see that there exists $\xi\in D(0,r)$ with $|(f^n)'(\xi)|>L$ and $|f^n(\xi)|= K$.
It follows that
\begin{equation}\label{17h}
(f^n)^\#(\xi)> \frac{L}{1+K^2}= \frac{K}{r(1+K^2)}  \log\frac{M^{n-m}(R,f)}{K}.
\end{equation}
We may assume here that $r< K/(1+K^2)$ so that the first term on the right side is greater than~$1$.
From this we can deduce that
\begin{equation}\label{17i}
(f^n)^\#(\xi)\geq \log M^{n-m}(R,f)
\end{equation}
for large~$n$.
\end{proof}

\section{Proof of Theorems~\ref{thm5} and~\ref{thm6}} \label{proofthm5}
\begin{proof}[Proof of Theorem~\ref{thm5}]
We may assume that $0$ is a periodic point. Since~\cite[p.~13]{Hayman1964}
the Nevanlinna characteristic $T(r,f)$ and the Ahlfors-Shimizu characteristic $T_0(r,f)$ satisfy
\[
\left|T(r,f)-T_0(r,f)-\log^+|f(0)|\right|\leq \frac12\log 2,
\]
this yields that
\begin{equation}\label{1a}
\left|T(r,f^n)-T_0(r,f^n)\right|\leq C
\end{equation}
for some constant $C$ independent of~$n$.

Choosing $r_1>1$ such that $E(f)\subset D(0,r_1)$ and $r_2>r_1$ we then have
\[
f^k(U)\supset D(0,r_2)\backslash D(0,r_1)
\]
for some $k\in\N$ by Lemma~\ref{excep-set}. Hence
\begin{equation}\label{1d}
S(U,f^{n+k})\geq S(D(0,r_2)\backslash D(0,r_1) ,f^n)= S(r_2,f^n)- S(r_1,f^n),
\end{equation}
with $S(r,f)=S(D(0,r),f)$.

We use the standard estimates
\[
\begin{aligned}
\frac12 S(\sqrt{r},f)\log r
& = \int_{\sqrt{r}}^r\frac{S(\sqrt{r},f)}{t}dt
\leq \int_{0}^r\frac{S(t,f)}{t}dt
\\ &
= T_0(r,f)
\leq S(r,f)\log r+T_0(1,f)
\end{aligned}
\]
which may also be written as
\begin{equation}\label{1b}
\frac{T_0(r,f)-T_0(1,f)}{\log r}
\leq S(r,f)
\leq \frac{T_0(r^2,f)}{\log r}.
\end{equation}
Now \eqref{1d} and \eqref{1b} give
\[
S(U,f^{n+k})\geq
\frac{T_0(r_2,f^n)-T_0(1,f^n)}{\log r_2}-
\frac{T_0(r_1^2,f^n)}{\log r_1}
\geq
\frac{T_0(r_2,f^n)}{\log r_2}-
2\frac{T_0(r_1^2,f^n)}{\log r_1}
\]
and thus
\[
S(U,f^{n+k})\geq
\frac{T(r_2,f^n)}{\log r_2}-
2\frac{T(r_1^2,f^n)}{\log r_1}
-\frac{2C}{\log r_1}
\]
by \eqref{1a}. With the standard estimate
\[
T(r,f)\leq \log^+M(r,f)\leq \frac{R+r}{R-r}T(R,f)
\]
this yields
\begin{equation}\label{1c}
S(U,f^{n+k})\geq
\frac13 \frac{\log M(\frac12 r_2,f^n)}{\log r_2}-
2\frac{\log^+ M(r_1^2,f^n)}{\log r_1}
-\frac{2C}{\log r_1}.
\end{equation}

Lemma~\ref{lem-fe} implies that $M(r_2/2,f^n)\geq M^n(r_2/4,f)$ if $r_2$ was chosen large enough.
Thus \eqref{1c} yields
\begin{equation}\label{1f}
\begin{aligned}
S(U,f^{n+k})
&\geq
\frac13 \frac{\log M^n(\frac14 r_2,f^n)}{\log r_2}-
2\frac{\log^+ M^n(r_1^2,f)}{\log r_1}
-\frac{2C}{\log r_1}
\\ &
\geq
\frac13 \frac{\log M^n(\frac14 r_2,f^n)}{\log r_2}-
3\frac{\log^+ M^n(r_1^2,f)}{\log r_1},
\end{aligned}
\end{equation}
provided $r_1$ is chosen large enough.

With $R_1=M(r_1^2,f)$ and $R_2= M(r_2/4,f)$ this takes the form
\begin{equation}\label{1g}
\begin{aligned}
S(U,f^{n+k})
&\geq
\frac13 \frac{\log M^{n-1}(R_2,f)}{\log R_2}\frac{\log R_2}{\log r_2}-
3\frac{\log^+ M^{n-1}(R_1,f)}{\log R_1}\frac{\log R_1}{\log r_1}.
\end{aligned}
\end{equation}
Since $f$ is transcendental,
\[
\frac{\log M(r,f)}{\log r} \to\infty
\]
as $r\to\infty$. This implies that we may choose $r_1$ and $r_2$ such that
\[
\frac{\log R_1}{\log r_1} \geq 1
\quad\text{and}\quad
\frac{\log R_2}{\log r_2}\geq 12 \frac{\log R_1}{\log r_1}.
\]
We may also assume that $M^{n-1}(R_2,f)\geq 1$ for all~$n$. Thus \eqref{1g} yields
\begin{equation}\label{1h}
\begin{aligned}
S(U,f^{n+k})
&\geq
\left(4 \frac{\log^+ M^{n-1}(R_2,f)}{\log R_2}-3\frac{\log^+ M^{n-1}(R_1,f)}{\log R_1}\right)
\frac{\log R_1}{\log r_1}
\end{aligned}
\end{equation}
if $r_1$ was chosen large enough.

Since $M(r,f)\geq r$ for large~$r$, Lemma~\ref{lem-fe3} yields that
there exists $R_0$ such that for $r\geq R_0$ and $c\geq 1$ we have
\begin{equation}\label{1h1}
\log M^n(r^c,f)\geq c \log M^n(r,f).
\end{equation}
This is equivalent to saying that
\[
\frac{\log M^n(r,f)}{\log r}
\]
is a non-decreasing function of $r$ for $r\geq R_0$.
Since we may assume that  $R_1\geq R_0$  this yields
\[
\frac{\log M^{n-1}(R_2,f)}{\log R_2} \geq \frac{\log M^{n-1}(R_1,f)}{\log R_1}.
\]
It now follows from \eqref{1h} that
\[
S(U,f^{n+k}) \geq \frac{\log M^{n-1}(R_2,f)}{\log R_2}
\]
if $r_2$ (and hence $R_2$) is large enough.

We now choose $l$ such that $M^l(R,f)\geq R^{\log R_2}$ and deduce that
\[
S(U,f^{n+k}) \geq \frac{\log M^{n-1-l}(M^l(R_2,f),f)}{\log R_2}
 \geq \frac{\log M^{n-1-l}(R^{\log R_2},f)}{\log R_2}.
\]
Using~\eqref{1h1} again this finally yields
\[
S(U,f^{n+k}) \geq \log M^{n-1-l}(R,f).
\]
The conclusion now follows with $m=l+k+1$.
\end{proof}

\begin{proof}[Proof of Theorem~\ref{thm6}]
Choose $R\geq e^4$  with
$M^n(R,f)\to\infty$ as $n\to\infty$
such that
$U\subset D\!\left(0,\sqrt{R}\right)$.
Then
\[
\begin{aligned}
S(U,f^n)&\leq S\!\left(\sqrt{R},f^n\right)\leq \frac{2T_0(R,f^n)}{\log R}
\leq  \frac12 \left(T(R,f^n)+C\right)
\\ &
\leq \frac12 \left(\log M(R,f^n)+C\right)\leq \log M^n(R,f)
\end{aligned}
\]
for large $n$ by~\eqref{1a} and~\eqref{1b}. Increasing $R$ if necessary we may achieve that
this holds for all $n\in\N$.
\end{proof}

\section{Proof of Theorem~\ref{thm1}} \label{proofthm1}
We will use the following result \cite[Theorem~1.2]{Bergweiler2005} already quoted in the introduction.
\begin{lemma}\label{multiplier}
Let $f$ be a transcendental entire function and let $p\in\N$, $p\geq 2$.
Then there exists a sequence  $(a_k)$ of fixed points of $f^p$
such that $(f^p)'(a_k)\to\infty$ as $k\to\infty$.
\end{lemma}
\begin{proof}[Proof of Theorem~\ref{thm1}]
We apply this lemma for $p=2$. We may assume that all $a_k$ are repelling fixed points of $f^2$.
Then there exist $r_k>0$ such that $f^2$ is univalent in the disk $D_k=D(a_k,r_k)$. Moreover, we may
assume that there exists an increasing sequence $(\lambda_k)$ tending to $\infty$  such that
\begin{equation}\label{2p1}
\left|(f^2)'(z)\right|\geq \lambda_k>1
\quad\text{for} \  z\in D_k
\end{equation}
and that there exists a domain $W_k$ satisfying $\overline{W_k}\subset D_k$ such that
$f^2\colon W_k\to D_k$ is univalent. This implies that for every $n_k\in\N$ there exists
a domain $V_k$ satisfying $\overline{V_k}\subset D_k$ such that $f^{2n_k}\colon V_k\to D_k$ is univalent.

We put $D_0=U$.  By the Ahlfors islands theorem (see~\cite[Section~5]{Hayman1964} or \cite{Bergweiler1998}),
for each $k\in\N$ there exist $m_k\in\N$ and a subdomain $U_{k-1}$ of $D_{k-1}$ such that
$f^{m_k}\colon U_{k-1}\to D_j$ is univalent for some $j\in\{k,k+1,k+2\}$. We may assume that this
holds for $j=k$ since otherwise we may restrict to a subsequence of $(a_k)$.
Thus $f^{m_k}\colon U_{k-1}\to D_k$ is univalent for $k\in\N$.

We conclude that for each $l\in\N$ there exists subdomain $X_l$ of $U_0$ such that
\begin{equation}\label{2p}
f^{m_l}\circ f^{2n_{l-1}}\circ f^{m_{l-1}}\circ \dots \circ f^{m_2}\circ f^{2n_{1}}\circ f^{m_{1}}
\colon X_l\to D_l
\end{equation}
is univalent, with $\overline{X_{l+1}}\subset X_l$. It follows that there exists
\begin{equation}\label{2q}
z\in\bigcap_{l=1}^\infty X_l =\bigcap_{l=1}^\infty \overline{X_l}.
\end{equation}
We show that we can achieve $\chi(f,z)=\infty$ by choosing the sequence $(n_k)$ rapidly increasing.

In order to do so we note that once the sequences $(m_l)$ and $(U_l)$ are fixed, there
are also sequences $(\alpha_l)$ and $(\beta_l)$ of positive numbers such that
\begin{equation}\label{20c}
\left|(f^k)'(\zeta)\right|\geq \alpha_l
\quad\text{and}\quad
\left|(f^k)(\zeta)\right|\leq \beta_l
\quad\text{for}\ 0\leq k\leq m_l
\ \text{and}\ \zeta\in U_{l-1}.
\end{equation}
Here, as usual, $f^0(\zeta)=\zeta$, so for $k=0$ the first inequality just means that $\alpha_l\leq 1$.

With the sequence $(n_k)$ still to be determined, we define sequences $(N_l)$ and $(M_l)$ by
\begin{equation}\label{20a}
N_l=\sum_{k=1}^l (2n_j+m_j) \quad\text{and}\quad M_l=N_l-2n_l=N_{l-1}+m_l
\end{equation}
so that $M_l<N_l<M_{l+1}$. We may choose $(n_k)$ such that
\begin{equation}\label{20a1}
n_l=\frac12 (N_l-M_{l}) \geq \frac14 (N_l+m_{l+1})+1=\frac14 M_{l+1} +1
\end{equation}
and
\begin{equation}\label{20a2}
\lambda_l^{n_l/2} \geq \frac{1+\beta_{l+1}^2}{\prod_{k=1}^{l+2} \alpha_k}
\geq
\frac{1+\beta_{l+1}^2}{\prod_{k=1}^{l+1} \alpha_k}
\end{equation}
for all~$l$.

Suppose first that $n\in\N$ is such that $N_l<n\leq  M_{l+1}=N_l+m_{l+1}$ for some $l\in\N$.
We deduce from~\eqref{2p1}, \eqref{20c}, \eqref{20a1} and~\eqref{20a2} that
\begin{equation}\label{20b}
\begin{aligned}
\left|(f^n)'(z)\right|
&=
\left|(f^{n-N_l}\circ f^{2n_{l}}\circ f^{m_{l}}\circ \dots \circ f^{m_2}\circ f^{2n_{1}}\circ f^{m_{1}})(z)\right|
\\ &
\geq
\prod_{k=1}^{l}  \lambda_k^{n_k}
\cdot
\prod_{k=1}^{l+1} \alpha_k
\geq
\lambda_l^{n_l} \cdot
\prod_{k=1}^{l+1} \alpha_k
\geq
(1+\beta_{l+1}^2) \lambda_l^{n_l/2}
\\ &
\geq
(1+\beta_{l+1}^2) \lambda_l^{M_{l+1}/8}
\geq
(1+\beta_{l+1}^2) \lambda_l^{n/8}
\end{aligned}
\end{equation}
and hence, using~\eqref{20c} again, that
\begin{equation}\label{20b1}
\left|(f^n)^\#(z)\right|\geq \frac{\left|(f^n)'(z)\right|}{1+\beta_{l+1}^2}\geq \lambda_l^{n/8}.
\end{equation}
Suppose next that $ M_{l+1}<n\leq N_{l+1}=M_{l+1}+2n_{l+1}$ for some $l\in\N$.
Using the same arguments as before we find that
\begin{equation}\label{20b2}
\begin{aligned}
\left|(f^n)'(z)\right|
&=
\left|(f^{n-M_{l+1}}\circ f^{m_{l+1}}\circ f^{2n_{l}}\circ \dots \circ f^{m_2}\circ f^{2n_{1}}\circ f^{m_{1}})(z)\right|
\\ &
\geq \lambda_{l+1}^{(n-M_{l+1})/2} \cdot \prod_{k=1}^{l}  \lambda_k^{n_k}
\cdot \prod_{k=1}^{l+1} \alpha_k
\geq \lambda_{l+1}^{(n-M_{l+1})/2}
(1+\beta_{l+1}^2) \lambda_l^{n_l/2}
\end{aligned}
\end{equation}
if $n-M_{l+1}$ is even while
\begin{equation}\label{20b3}
\begin{aligned}
\left|(f^n)'(z)\right|
&=
\left|(f^{n-M_{l+1}}\circ f^{m_{l+1}}\circ f^{2n_{l}}\circ \dots \circ f^{m_2}\circ f^{2n_{1}}\circ f^{m_{1}})(z)\right|
\\ &
\geq \alpha_{l+2}\cdot \lambda_{l+1}^{(n-M_{l+1}-1)/2} \cdot \prod_{k=1}^{l}  \lambda_k^{n_k}
\cdot \prod_{k=1}^{l+1} \alpha_k
\geq \lambda_{l+1}^{(n-M_{l+1}-1)/2}
(1+\beta_{l+1}^2) \lambda_l^{n_l/2}
\end{aligned}
\end{equation}
if $n-M_{l+1}$ is odd. Thus
\begin{equation}\label{20d}
\left|(f^n)'(z)\right|\geq \lambda_{l+1}^{(n-M_{l+1}-1)/2} (1+\beta_{l+1}^2) \lambda_l^{n_l/2}
\end{equation}
in both cases. Since $(\lambda_k)$ is increasing and
\begin{equation}\label{20d1}
n-M_{l+1}-1+n_l\geq n-\frac34 M_{l+1} \geq \frac{n}{4}
\end{equation}
by~\eqref{20a1}, we find that
\begin{equation}\label{20e}
\left|(f^n)'(z)\right|\geq(1+\beta_{l+1}^2) \lambda_{l}^{(n-M_{l+1}-1+n_l)/2} \geq (1+\beta_{l+1}^2)\lambda_l^{n/8},
\end{equation}
which is the same inequality as~\eqref{20b}.  We conclude that~\eqref{20b1} and hence
\begin{equation}\label{20f}
\frac{1}{n}\log\left|(f^n)^\#(z)\right|\geq \frac{1}{8} \log\lambda_l
\end{equation}
holds for all $n\geq N_1$.
Since $l$ and hence $\lambda_l$ tend to $\infty$ with~$n$, this yields $\chi(f,z)=\infty$.

To prove that the set of all $\zeta$ with $\chi(f,\zeta)=\infty$ is dense in $J(f)$ we note that
if this holds for $\zeta=z$, then it also holds for $\zeta=f^n(z)$ if $n\in\N$. More
generally, it holds  for all $\zeta$ for which there exist $m,n\in\N$ such that
$f^m(\zeta)=f^n(z)$ and $(f^m)'(\zeta)\neq 0$. The set of all such $\zeta$ is easily seen
to be dense in $J(f)$, using the Ahlfors island theorem -- or the simpler result that if
$a_1,a_2,a_3\in\C$ are distinct, then the family of all functions holomorphic in a domain
which have no simple $a_j$-points for all $j$ is normal.
\end{proof}
\begin{remark} \label{rem-dense}
Given a sequence $(V_k)$ of open sets intersecting $J(f)$, one may choose the sequences
$(m_k)$ and $(U_k)$ in the above proof such that $f^{l_k}(U_{k-1})\subset V_k$ for some $l_k\leq m_k$.
Using this it is not difficult to see that one may choose $z$ with the additional property
that the orbit of $z$ is dense in $J(f)$.

Similarly, given any sequence $(c_k)$ of positive real numbers tending to~$\infty$, one may choose
$z$ such that $|f^k(z)|\leq c_k$ for all large~$k$.
This can be achieved by choosing $(n_k)$ large; a similar idea appears in~\cite{Rippon2011}.
\end{remark}

\section{Proof of Theorem~\ref{thm3}} \label{proofthm2}
We recall the logarithmic change of variable for a function
$f$ in the Eremenko-Lyubich class; see \cite[\S 2]{Eremenko1992}. For simplicity we will assume that
all singularities of the inverse are in the unit disk and that $|f(0)|<1$.
The general case can be reduced to this.
Let  $U$ be a logarithmic tract of~$f$, that is, a component of $\{z\in\C\colon |f(z)|>1\}$.
Let $H=\{z\in\C\colon\re z>0\}$ be the right half-plane and $W=\exp^{-1}(U)$.
Then there exists a $2\pi i$-periodic holomorphic function
$F\colon W\to H$ satisfying $\exp F(z)=f(e^z)$, and the restriction of $F$ to
a component of $W$ maps this component biholomorphically onto~$H$.

We call $F$ the function obtained from $f$ by a logarithmic change of variable.
The main tool when working with the Eremenko-Lyubich class is the inequality
\begin{equation}\label{4a0}
|F'(z)|\geq\frac{1}{4\pi}\re F(z)
\quad\text{for}\ z\in W
\end{equation}
obtained by them. We will also need the following lower bound for~$F'$.
\begin{lemma}\label{lemma1a}
Let $F\colon W\to H$ be  the function obtained from a logarithmic change of variable as above.
For $z\in W$ let $z_1\in \partial W$ with $|z_1-z|=\dist(z,\partial W)$. Then
\begin{equation}\label{9a}
|F'(\zeta)|\geq\frac{1}{8\pi}\re F(z)
\end{equation}
for all $\zeta$ in the straight line segment from $z$ to $z_1$.
\end{lemma}
\begin{proof}
Let $G\colon H\to W$ be the branch of the inverse of $F$ with $G(F(\zeta))=\zeta$.
Since $G$ is univalent in $D(F(\zeta),\re F(\zeta))$, Koebe's one quarter theorem yields that
\begin{equation}\label{9a1}
W\supset G(H)\supset G(D(F(\zeta),\re F(\zeta)))\supset D\!\left(\zeta,\frac14 |G'(F(\zeta))|\re F(\zeta)\right)
\end{equation}
and hence
\begin{equation}\label{9a2}
\frac{\re F(\zeta)}{4 |F'(\zeta)|}=
\frac14 |G'(F(\zeta))|\re F(\zeta)\leq \dist(\zeta,\partial W)=|\zeta-z_1|.
\end{equation}
We note that \eqref{4a0} follows from this by noting that $\dist(\zeta,\partial W)\leq\pi$.
In fact, this is the proof of~\eqref{4a0} given in~\cite{Eremenko1992}.

To prove~\eqref{9a}, we write $\zeta=z+s(z_1-z)=sz_1+(1-s)z$ with $0<s<1$  and put $u(z)=\re F(z)$.
Harnack's inequality yields that
\begin{equation}\label{9b}
u(\zeta)\geq \frac{1-s}{1+s}u(z).
\end{equation}
Hence
\begin{equation}\label{9c}
|\zeta-z_1|=(1-s) |z-z_1| \leq 2\frac{u(\zeta)}{u(z)}|z-z_1|
=2|z-z_1|\frac{\re F(\zeta)}{\re F(z)}.
\end{equation}
Together with~\eqref{9a2} this yields
\begin{equation}\label{9d}
\frac{1}{4 |F'(\zeta)|}\leq \frac{2|z-z_1|}{\re F(z)},
\end{equation}
from which the conclusion follows since $|z-z_1|\leq\pi$.
\end{proof}

For $f\in B$ and the function $F\colon W\to H$ obtained
from $f$ by the logarithmic change of variable we
put $\alpha=\inf\{\re z\colon z\in W\}$. As in~\cite[\S 3]{Bergweiler2009} we consider the function
$h\colon(\alpha,\infty)\to (0,\infty)$ defined by
\[
h(x)=\max_{\re z=x} \re F(z) =\max_{y\in\R} \re F(x+iy).
\]
Note that $h$ is increasing by the maximum principle. Moreover,
$h$ is convex by analogy to Hadamard's three circles theorem.
\begin{lemma}\label{lemma1}
For $x>\alpha$ let $z_x\in W$ with
\begin{equation}\label{5a0}
\re z_x=x
\quad\text{and}\quad
\re F(z_x)=h(x).
\end{equation}
Then for each $t\in (0,h(x)]$ there exists $\zeta_t \in D(z_x,\dist(z_x,\partial W))$ such that
\begin{equation}\label{4h1}
\re F(\zeta_t)=t \quad\text{and}\quad |F'(\zeta_t)|\geq \frac{h(x)}{8\pi}.
\end{equation}

Moreover, if $t\geq 8\pi$ and if $U_t$ is the component of
$F^{-1}(D(F(\zeta_t),4\pi))$ that contains~$\zeta_t$, then
\begin{equation}\label{4h0}
|F'(z)| \geq  \frac{h(x)}{96\pi} \quad\text{for}\ z\in U_t
\end{equation}
and
\begin{equation}\label{4j0}
U_t\subset D(\zeta_t,2).
\end{equation}
\end{lemma}
\begin{proof}
It follows from Lemma~\ref{lemma1a} that there exists $\zeta_t\in D(z_x,\dist(z_x,\partial W))$
satisfying~\eqref{4h1}.

Let now $t\geq 8\pi$ and, as in the proof of Lemma~\ref{lemma1a},
let $G\colon H\to W$ be the branch of the inverse of $F$ with $G(F(\zeta_t))=\zeta_t$.
Since $G$ is univalent in $D(F(\zeta_t),\re F(\zeta_t))$, Koebe's distortion theorem implies that
\begin{equation}\label{4g}
|G'(w)| \leq |G'(F(\zeta_t))|\frac{1+4\pi/t}{(1-4\pi/t)^3}\leq 12 |G'(F(\zeta_t))|
\quad\text{for}\ w\in D(F(\zeta_t),4\pi)
\end{equation}
and hence that
\begin{equation}\label{4h}
|F'(z)| \geq  \frac{1}{12}|F'(\zeta_t)|
\quad\text{for}\ z\in U_t ,
\end{equation}
which together with~\eqref{4h1} yields~\eqref{4h0}.
Koebe's distortion theorem and~\eqref{4h1} also yield that if $z\in U_t$, then
\begin{equation}\label{4i}
|z-\zeta_t|
\leq |G'(F(\zeta_t))| \frac{4\pi/t}{(1-4\pi/t)^2}
\leq  |G'(F(\zeta_t))| \frac{16\pi}{t}
=  \frac{16\pi}{|F'(\zeta_t)|t} \leq  \frac{128\pi^2}{h(x)t}.
\end{equation}
Since $8\pi\leq t\leq h(x)$ this yields~\eqref{4j0}.
\end{proof}

\begin{lemma}\label{lemma2}
Let $(x_n)_{n\geq 0}$ be a sequence of positive numbers satisfying
\begin{equation}\label{5a1}
x_n> \max\{\alpha,8\pi\}
\quad\text{and}\quad
x_{n+1}\leq h(x_n)
\end{equation}
for all $n\geq 0$. Then there exists $u\in W$ such that
\begin{equation}\label{5a}
\left|\re F^n(u)-x_n\right|\leq 4\pi
\quad\text{and}\quad
\left|(F^n)'(u)\right|\geq \frac{1}{(96\pi)^n} \prod_{j=0}^{n-1} h(x_j)
\end{equation}
for all $n\geq 1$.
\end{lemma}
\begin{proof}
First we choose $z_0$ with $\re z_0=x_0$ and $\re F(z_0)=h(x_0)$.
By Lemma~\ref{lemma1} there exist $\zeta_0\in D(z_0,2)$ such that
with $\xi_1= F(\zeta_0)$ we have
$\re \xi_1=x_1$ and such that the component $V_1$ of $F^{-1}(D(\xi_1,4\pi)$
that contains $\zeta_0$ satisfies $V_1\subset D(\zeta_0,2)$ and
\begin{equation}\label{5b}
|F'(z)|\geq \frac{1}{96\pi} h(x_0)
\quad\text{for}\ z\in V_1 .
\end{equation}
We now choose a point $z_1$ with $\re z_1=\re \xi_1=x_1$ and $|\im z_1-\im \xi_1|\leq\pi$ such that $\re F(z_1)=h(x_1)$.
Using Lemma~\ref{lemma1} again, we see that there exists $\zeta_1\in D(z_1,\pi)\subset D(\xi_1,2\pi)$ such that
with $\xi_2= F(\zeta_1)$ we have $\re \xi_2=x_2$, and the component $U_2$ of $F^{-1}(D(\xi_2,4\pi)$
that contains $\zeta_1$ satisfies $U_2\subset D(\zeta_1,2)$ and
\begin{equation}\label{5c}
|F'(z)|\geq \frac{1}{96\pi} h(x_1)
\quad\text{for}\ z\in U_2.
\end{equation}
Note that that $U_2\subset D(\xi_1,2+2\pi)$.
Since $F\colon V_1\to D(\xi_1,4\pi)$ is biholomorphic we deduce that
there exists a domain $V_2$ satisfying $\overline{V_2}\subset V_1$ such
that $F\colon V_2\to U_2$ is biholomorphic.
Hence $F^2\colon V_2\to D(\xi_2,4\pi)$ is biholomorphic. Moreover, it follows from~\eqref{5b} and~\eqref{5c} that
\begin{equation}\label{5d}
|(F^2)'(z)|\geq \frac{1}{(96\pi)^2} h(x_0)h(x_1)
\quad\text{for}\ z\in V_2.
\end{equation}
Inductively we thus find a sequence $(\xi_n)$ of points satisfying
$\re \xi_n=x_n$ and a sequence $(V_n)$ of domains satisfying $\overline{V_n}\subset V_{n-1}$ such
that $F^n\colon V_{n}\to D(\xi_n,4\pi)$ is biholomorphic
\begin{equation}\label{5e}
|(F^n)'(z)|\geq \frac{1}{(96\pi)^n} \prod_{j=0}^{n-1}h(x_j)
\quad\text{for}\ z\in V_n.
\end{equation}
The conclusion now follows by choosing $u\in\bigcap_{n=1}^\infty V_n$.
\end{proof}

\begin{proof}[Proof of Theorem~\ref{thm3}]
By hypothesis we have
\begin{equation}\label{5f}
\log M(r,f)\geq r^{\lambda(f)-o(1)}
\end{equation}
as $r\to\infty$. In terms of $F$ this takes the form
\begin{equation}\label{6a}
h(x)\geq e^{(\lambda-\varepsilon(x))x}
\end{equation}
where $\lambda=\lambda(f)$ and $\varepsilon(x)\to 0$.
We may assume here that $\varepsilon$ is non-increasing, since otherwise we may replace it by
\begin{equation}\label{6a1}
\varepsilon^*(x)=\sup_{t\geq x} \varepsilon(t).
\end{equation}
We now consider, for $x>1$,
\begin{equation}\label{6a2}
\delta(x)=\max\left\{\varepsilon(x),\frac{1}{\log x}\right\}.
\end{equation}
Then $\delta$ is non-increasing. We now choose $x_0$ large,
\begin{equation}\label{6a3}
x_{1}=\frac{\lambda-\delta(x_0)}{1+\delta(x_0)}x_0
\quad\text{and}\quad
x_{n+1}=\frac{1+\lambda}{1+\delta(x_n)}x_n
\end{equation}
for $n\geq 1$.
It follows from~\eqref{6a3} that there exists a sequence $(\eta_n)$ tending to $0$ such that
\begin{equation}\label{6f}
x_n= ( 1+\lambda+\eta_n)^n,
\end{equation}
provided $x_0$ was chosen large enough.

Induction shows that
\begin{equation}\label{6b}
\sum_{j=0}^{n-1}(\lambda-\delta(x_j))x_j\geq (1+\delta(x_{n-1}))x_n
\end{equation}
for all $n\geq 1$. Indeed, this holds for $n=1$ by the choice of $x_1$ and assuming
that~\eqref{6b} holds we obtain, using that $\delta(x)$ is non-increasing,
\begin{equation}\label{6c}
\begin{aligned}
\sum_{j=0}^{n}(\lambda-\delta(x_j))x_j
&
=\sum_{j=0}^{n-1}(\lambda-\delta(x_j))x_j + (\lambda-\delta(x_n)) x_n
\\ &
\geq (1+\delta(x_{n-1}))x_n+ (\lambda-\delta(x_n)) x_n
\\ &
= (1+\lambda+\delta(x_{n-1})-\delta(x_n)) x_n
\\ &
\geq
(1+\lambda)x_n
= (1+\delta(x_n))x_{n+1}.
\end{aligned}
\end{equation}
It follows from~\eqref{6f} that Lemma~\ref{lemma2} is applicable if $x_0$ was chosen large enough.
With $u$ as in this lemma we thus have
\begin{equation}\label{6d}
\begin{aligned}
\left|(F^n)'(u)\right|
&\geq \frac{1}{(96\pi)^n} \prod_{j=0}^{n-1} h(x_j)
\\ &
\geq \exp\!\left(\sum_{j=0}^{n-1} (\lambda-\delta(x_j))x_j - n\log\!\left(96\pi\right)\right)
\\ &
\geq \exp\!\left((1+\delta(x_{n-1}))x_n - n\log\!\left(96\pi\right)\right).
\end{aligned}
\end{equation}
Since $\exp F(u)=f(e^u)$ we find, with $z=e^u$, that
\begin{equation}\label{5m}
(F^n)'(u)=z\frac{(f^n)'(z)}{f^n(z)}.
\end{equation}
Since $\re F^n(u)\to\infty$ and thus $|f^n(z)|\to\infty$ by~\eqref{6f} and Lemma~\ref{lemma2}, we have
\begin{equation}\label{5n}
(f^n)^\#(z)\geq \frac{|(f^n)'(z)|}{2|f^n(z)^2|}=\frac{|(F^n)'(u)|}{2|z|\exp(\re F^n(u))}
\end{equation}
for large~$n$. Combined with~\eqref{6a2} and~\eqref{6d} this yields
\begin{equation}\label{6e}
\begin{aligned}
(f^n)^\#(z)
&\geq \exp\!\left((1+\delta(x_{n-1}))x_n - n\log (96\pi)- x_n-4\pi \log(2|z|) \right)
\\ &
= \exp\!\left(\delta(x_{n-1})x_n - n\log (96\pi)-4\pi \log(2|z|) \right)
\\ &
\geq \exp\!\left(\frac{x_n}{\log x_n} - n\log (96\pi)-4\pi \log(2|z|) \right).
\end{aligned}
\end{equation}
For large $n$ we thus have
\begin{equation}\label{6g}
(f^n)^\#(z) \geq \exp\!\left(\frac{x_n}{2 \log x_n} \right)
\end{equation}
and hence
\begin{equation}\label{6h}
\begin{aligned}
\log\log  (f^n)^\#(z)
&\geq \log x_n -\log\log x_n -\log 2
\\ &
= n \log( 1+\lambda+\eta_n) - \log n- \log \log( 1+\lambda+\eta_n) -\log 2,
\end{aligned}
\end{equation}
from which \eqref{5p} follows.

Once it is known that there exists one point $z$ satisfying~\eqref{5p}, it follows as in the proof
of Theorem~\ref{thm1} that the set of all such $z$ is dense in $J(f)$.
\end{proof}

\section{Proof of Theorem~\ref{thm4}} \label{proofthm4}
Mittag-Leffler's function
\begin{equation}\label{7a}
E_\alpha(z)=\sum^\infty_{n=0} \frac{z^n}{\Gamma(\alpha n +1)}
\end{equation}
satisfies $\rho(E_\alpha)=\lambda(E_\alpha)=1/\alpha$.
It was shown in~\cite[Section~4]{Aspenberg2012} that $f$ is in the Eremenko-Lyubich class if $0<\alpha<2$.
Since $E_{2}(z)=\cosh\sqrt{z}$ this also holds for $\alpha=2$.

For $0<\alpha<2$ and $\rho=1/\alpha$ we have (see \cite[p.~ 85]{Goldberg2008}
and~\cite[Section~4]{Aspenberg2012}), for sufficiently small $\delta>0$,
\begin{align}
E_{\alpha}(z) &= \varrho \exp\left(z^{\varrho}\right) + O\!\left(\frac{1}{|z|}\right)
\quad \text{for} \ |\arg (z)| \leq \frac{\alpha\pi}{2} +\delta, \label{7b} \\
E_{\alpha}'(z) &= \varrho^2 z^{\varrho-1} \exp\left(z^{\varrho}\right) + O\!\left(\frac{1}{|z|^2}\right)
\quad \text{for} \ |\arg (z)| \leq \frac{\alpha\pi}{2}+\delta , \label{7c} \\
E_{\alpha}(z) &= O\!\left(\frac{1}{|z|}\right)
\quad \text{for} \ \frac{\alpha\pi}{2}+\delta < | \arg (z) | \leq \pi,  \label{7d} \\
E_{\alpha}'(z) &= O\!\left(\frac{1}{|z|^2}\right)
\quad \text{for} \ \frac{\alpha\pi}{2}+\delta < | \arg (z) | \leq \pi.  \label{7e}
\end{align}
This implies that there exists constants $A$ and $B$ such that
\begin{equation}\label{7f}
|E_\alpha'(z)|\leq A |z|^{\rho-1} |E_\alpha(z)| +B
\end{equation}
for all $z\in\C$. With $C=A+B$ we thus have
\begin{equation}\label{7i}
|E_\alpha'(z)|\leq C |z|^{\rho-1} |E_\alpha(z)|
\quad\text{if}\ |z|\geq 1 \ \text{and}\ |E_\alpha(z)| \geq 1.
\end{equation}
Since $E_{2}(z)=\cosh\sqrt{z}$ the last estimate also holds for $\alpha=2$.

We consider the function
\begin{equation}\label{7g}
f(z)= \eta E_\alpha(z)
\end{equation}
where $0<\eta<1$. Since $E_\alpha\in B$ we have $f\in B$.
By choosing $\eta$ small we can achieve that
\begin{equation}\label{7h}
|f(z)|<1 \quad\text{and}\quad |f'(z)|<1 \quad\text{for}\ |z|\leq 1.
\end{equation}
Moreover, \eqref{7i} implies that
\begin{equation}\label{7i1}
|f'(z)|\leq C |z|^{\rho-1} |f(z)|
\quad\text{if}\ |z|\geq 1 \ \text{and}\ |f(z)| \geq 1.
\end{equation}

Suppose now that $z$ satisfies $\chi(f,z)=\infty$ and put $z_n=f^n(z)$ for $n\geq 0$ so that $z_0=z$.
It follows from~\eqref{7h} that if $|z_N|\leq 1$ for some $N\geq 0$, then
$(f^n)'(z)\to 0$ and hence $(f^n)^\#(z)\to 0$ as $n\to\infty$.
Thus we may assume that $|z_n|\geq 1$ for all $n\geq 0$. Using~\eqref{7i1} we see that
\begin{equation}\label{7j}
|(f^n)'(z)| =\prod_{j=0}^{n-1} \left|f'(z_{j})\right|
\leq C^n\prod_{j=0}^{n-1} |z_{j}|^{\rho-1} |z_{j}|
= \frac{C^n|z_n|}{|z_0|} \prod_{j=0}^{n-1} |z_{j}|^{\rho} .
\end{equation}
Hence
\begin{equation}\label{7k}
(f^n)^\#(z)\leq \frac{|(f^n)'(z)|}{|(f^n)(z)|^2}\leq \frac{C^n}{|z_n|} \prod_{j=0}^{n-1} |z_{j}|^{\rho} .
\end{equation}
If
\begin{equation}\label{7l}
|z_n|> \prod_{j=0}^{n-1} |z_{j}|^{\rho} ,
\end{equation}
then~\eqref{7k} yields that $(f^n)^\#(z)\leq C^n$. Since $\chi(f,z)=\infty$, we deduce that~\eqref{7l}
cannot hold for infinitely many~$n$. Thus there exists $n_0\in\N$ such that
\begin{equation}\label{7n}
|z_n|\leq  \prod_{j=0}^{n-1} |z_{j}|^{\rho}
\quad\text{for}\ n\geq n_0.
\end{equation}
We put $t_n=\log|z_n|$. Then the last inequality takes the form
\begin{equation}\label{7o}
t_n\leq \rho \sum_{j=0}^{n-1} t_{j} \quad\text{for}\ n\geq n_0.
\end{equation}
This implies that
\begin{equation}\label{7p}
\sum_{j=0}^{n} t_{j}= t_n + \sum_{j=0}^{n-1} t_{j} \leq (1+\rho) \sum_{j=0}^{n-1} t_{j}.
\end{equation}
Induction yields that
\begin{equation}\label{7q}
\sum_{j=0}^{n} t_{j} \leq c_0(1+\rho)^n \quad\text{for}\ n\geq n_0-1,
\end{equation}
with
\begin{equation}\label{7r}
c_0= (1+\rho)^{-n_0+1} \sum_{j=0}^{n_0-1} t_{j}.
\end{equation}
Using~\eqref{7k} and~\eqref{7q} we find for $n\geq n_0$ that
\begin{equation}\label{7s}
\begin{aligned}
\log (f^n)^\#(z)
&\leq \log\!\left( C^n \prod_{j=0}^{n-1} |z_{j}|^{\rho} \right)
 = n\log C + \rho \sum_{j=0}^{n-1} t_{j}
\leq n\log C + \rho c_0(1+\rho)^{n-1}.
\end{aligned}
\end{equation}
Hence
\begin{equation}\label{7t}
\log (f^n)^\#(z) =O\!\left( (1+\rho)^{n} \right)
\end{equation}
as $n\to\infty$, which yields~\eqref{5p3}.

\begin{remark} \label{rem-IA}
The proof of Theorem~\ref{thm3} shows that if $f\in B$ or, more generally, if $f$ has a logarithmic
singularity over~$\infty$, then there exists $z\in I(f)$ satisfying~\eqref{5p}. In particular,
there exists $z\in I(f)$ with $\chi(f,z)=\infty$.
We do not know whether this holds for all transcendental entire functions~$f$.

On the other hand, it follows from the proof of Theorem~\ref{thm4} that in general there does
not exist $z\in A(f)$ with $\chi(f,z)=\infty$.  Indeed, it is easily seen that if $f$ is as there,
$z\in A(f)$ and $z_n=f^n(z_n)$, then~\eqref{7l} holds for large~$n$. As shown in the proof
of  Theorem~\ref{thm4} this is incompatible with $\chi(f,z)=\infty$.
\end{remark}

\end{document}